I have a story problem that is giving me a problem could somebody please help?
Mr.Adams buys a bag of apples and a bag of pears for a total of 14$. There are eight pounds of fruit in all. The apples cost $1.50 per pound, and the pears cost $2.30 perpound. HOw many pounds of pears did he buy?

Answer by Paul(988)

Let x represent the number of apples
Let y represent the number of pears
Total = 8

x+y=8
y=8-x

Cost of apples = $1.50
Cost of pears = $2.30
Total cost = $14

1.50x+2.30y=14 (subsitute)
1.50x+2.30(8-x)=14
1.50x+18.40-2.30x=14
-0.8x =-4.4 Divide both sides by -0.8
x = 5.5

y=8-5.5
y=2.5

Hence there were 2.5 pounds of pears and 5.5 pounds of apples.
Check. $1.50(5.5)+$2.3(2.5)=$14
Paul.